# Edwards-Wilkinson limit for a stochastic advection-diffusion PDE

### v1

@@ Abstract.
We consider a diffusion in a Gaussian random environment that is white in time, and study the large-scale behavior of the quenched density with respect to the Lebesgue measure. We show that under diffusive rescaling, the fluctuations of the density converge to a Gaussian limit, described by an additive stochastic heat equation. In the case where the environment is divergence-free, our result can be interpreted as computing the scaling limit of the first-order correction to the quenched Central Limit Theorem.
@@ 1. Introduction
for an appropriate function $Q:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$.
@@
(1.3) $\mathrm{d}X_{t}=V(t,X_{t})\,\mathrm{d}t+\sqrt{2\kappa}\,\mathrm{d}B_{t}.$
Then, for almost every realization of $V$, $\theta(t,x)$ is the density of $X_{t}$ with respect to the Lebesque measure.
@@
As such, studying the large-scale behavior of (1.1), yields ‘local’ information for a diffusion in a (white-in-time) random environment. This can be done in different scaling regimes. Specifically, [1] distinguishes three different scaling regimes: the diffusive regime, the moderate deviation regime, and the large deviation regime. These correspond to studying the diffusion (1.3) under a diffusive rescaling, under a moderate tilting of the diffusion, or the large deviation behavior, respectively. Moreover, in [1], the authors point out a very interesting connection to the KPZ equation and the KPZ universality class [3]. These conjectures were recently proved rigorously in [24, 8] in $d=1$, and for the moderate deviation regime [^1] The model studied in [24, 8] is a discrete analog of (1.1). However, their methods also apply to the continuous case, as is pointed out in [8], Section 6.3. We also refer to [4] for a study of a related, integrable model, under the large deviation regime.
Here, we are interested in the behavior of (1.1) in the diffusive scaling regime. In [1], it is conjectured that, in this regime, the fluctuations of $\theta(t,x)$, viewed as a random field, fall into the Edwards-Wilkinson universality class. To this end, we point out again the reference [10]. There, the authors studied the point-wise behavior of $\theta(t,x)$ and proved that
@@
as $n\rightarrow\infty$, where $q_{t}(x)$ is the standard $d$-dimensional heat kernel with a specific diffusivity and $\Psi(t,x)$ is an appropriate space-time stationary random field (see [10, Section 3] for more details). Instead, what we are interested in is the behavior of
(1.4) $\mathcal{X}_{n}(t,x):=n^{d/2}(\theta_{n}(t,x)-\mathbb{E}[\theta_{n}(t,x)]),$
where $\theta_{n}(t,x):=n^{d}\theta(n^{2}t,nx)$, viewed as a random element of a Sobolev space with negative order. Our main result shows that $\mathcal{X}_{n}(t,x)$ converges in distribution to an explicit Gaussian limit, confirming the predictions of [1], see Theorem 1.9 for the precise statement.
@@
We end this introduction by describing the structure of the rest of the paper. In Section 1.2 we set up the solution theory for (1.1) and state our assumptions, then we present our main result in Section 1.3 and briefly describe our methods. In Section 2.1 we collect estimates on the correlation functions of the model and prove some a priori estimates for (1.1). In Section 2.2 we prove our main result, while in Section 2.3 we show that our result still holds under weaker assumptions on the correlation function of the noise, if we assume that the latter is divergence-free.
@@ 1.1. Notation
For $x\in\mathbb{R}^{d}$, we write $x^{T}$ or $x^{\ast}$ as the transposition of $x$. The notation $|x|$ stands for the usual Euclidean norm while $\langle x\rangle=(1+|x|^{2})^{1/2}$. Let $\mathcal{S}$ be the space of Schwartz test functions and denote the Fourier transform of a function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$ as
@@
Given $T>0$ and $p,q\geq 1$, we denote $L^{p}_{t}L^{q}_{x}$ for the time-dependent space $L^{p}([0,T],L^{q}(\mathbb{R}^{d}))$; similarly, $C_{t}L^{q}_{x}$ and $C_{t}H^{\alpha}_{x}$ are abbreviations of $C([0,T],L^{q}(\mathbb{R}^{d}))$ and $C([0,T],H^{\alpha}(\mathbb{R}^{d}))$, respectively. Sometimes, we replace the subscript $t$ by $T$ to stress the length of the time interval $[0,T]$. We write $a\lesssim b$ to mean that there is some unimportant constant $C>0$ such that $a\leqslant Cb$; to emphasize the dependence of $C$ on some parameters $d,\kappa$, we use the notation $a\lesssim_{d,\kappa}b$.
Further, we will denote by $q_{t}(x)=\frac{1}{(2\pi t)^{d/2}}e^{-\frac{|x|^{2}}{2t}}$ the standard $d$-dimensional heat kernel.
Finally, we will make use of the notation $x_{1:p}=(x_{1},\dots,x_{p})$ and for a function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we define $g^{\otimes p}:\mathbb{R}^{pd}\rightarrow\mathbb{R}$, $g^{\otimes p}(x_{1:p}):=g(x_{1})g(x_{2})\dots g(x_{p})$.
@@ 1.2. The Setup and Assumptions
(1.5) $\widehat{Q}(\xi)=g(\xi)\bigg{(}I_{d\times d}-\frac{\xi\xi^{T}}{|\xi|^{2}}\bigg{)},$
where $g(\xi)=g(|\xi|)$ is a nonnegative radial function satisfying $g\in(L^{1}\cap L^{\infty})(\mathbb{R}^{d})$. It is easy to show that $Q(0)=2\nu I_{d\times d}$ for some $\nu>0$.
Observe that the matrix appearing in the right-hand side of (1.5) is the projection to the subspace orthogonal to $\xi$, so that $Q$ is indeed divergence-free.
@@
The matrix $Q$ is smooth and compactly supported, such that $Q(0)=2\nu I_{d\times d}$.
@@
It can be proved that the space $\mathcal{H}$ consists of continuous, bounded vector fields. We refer to [16, Lemma 2.2] and the discussion below for more details. We also record the following lemma from the same paper, see Lemma $2.3$ therein [^3] Strictly speaking, in [16] the authors consider only the case where Assumption 1.1 holds, but it is easy to see that similar arguments can be used to establish the same results under Assumption 1.2.:
@@
$Q(x-y)=\sum_{n\in\mathbb{N}}\sigma_{k}(x)\sigma_{k}(y)^{T},$
@@
$\sum_{k\in\mathbb{N}}\widehat{\sigma}_{k}(\xi)\overline{\widehat{\sigma}_{k}(\eta)^{T}}=\widehat{Q}(\xi)\delta(\xi-\eta),\quad\xi,\eta\in\mathbb{R}^{d},$
where the overline means complex conjugate, and $\delta(\xi-\eta)=1$ if $\xi=\eta$ and 0 otherwise.
@@
$\lim_{N\to\infty}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\phi(\xi)^{T}\Bigg{(}\sum_{k=0}^{N}\widehat{\sigma}_{k}(\xi)\overline{\widehat{\sigma}_{k}(\eta)^{T}}\Bigg{)}\overline{\psi(\eta)}~\,\mathrm{d}\xi\,\mathrm{d}\eta$
$=\lim_{N\to\infty}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\phi}(x)^{T}\Bigg{(}\sum_{k=0}^{N}\sigma_{k}(x)\sigma_{k}(y)^{T}\Bigg{)}\overline{\widehat{\psi}(y)}~\,\mathrm{d}x\,\mathrm{d}y$
$=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\phi}(x)^{T}\,Q(x-y)\,\overline{\widehat{\psi}(y)}~\,\mathrm{d}x\,\mathrm{d}y$
$=\int_{\mathbb{R}^{d}}\phi(\xi)^{T}\widehat{Q}(\xi)\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\overline{\operatorname{e}^{{\rm i}y\cdot\xi}\widehat{\psi}(y)}~\,\mathrm{d}y\,\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}\phi(\xi)^{T}\widehat{Q}(\xi)\overline{\psi(\xi)}~\,\mathrm{d}\xi,$
where in the second identity we used the fact that $\sum_{k=0}^{N}\sigma_{k}(x)\sigma_{k}(y)^{T}\to Q(x-y)$ uniformly on any compact sets. ∎
@@
and can be written as Itô integral plus correction, that is
@@
where $\nu$ is as in Assumptions 1.1, 1.2. This leads us to the following notion of solution to (1.1) (see also [16, Definition 2.16]):
@@
Let $(\Omega,\mathcal{A},(\mathcal{F})_{t},\mathbb{P})$ be a given filtered probability space satisfying the standard assumptions, let $V$ be as above. Let $\theta_{0}\in L^{1}\cap L^{p}$, for some $p\in(1,\infty)$. A solution to (1.1) is an $(\mathcal{F})_{t}$-progressively measurable process $\theta:[0,T]\times\Omega\to L^{1}\cap L^{p}$, satisfying
@@
    $\int_{0}^{t}\langle\theta_{s},\nabla\phi\cdot V(\,\mathrm{d}s)\rangle=\sum_{k}\int_{0}^{t}\langle\nabla\phi,\sigma_{k}\theta_{s}\rangle\,\mathrm{d}B^{k}_{s}.$
@@
It can be shown, see for example [16, Appendix B], that, when the spectral intensity $g$ in (1.5) is decaying rapidly at infinity, that is, when $Q$ is sufficiently smooth, the Stratonovich formulation (1.1) makes sense and the equivalence with the Itô formulation (1.7) holds rigorously. Hence, the point (ii) in Definition 1.5 is equivalent to:
@@
In the case $Q$ satisfies Assumption 1.2, if the initial condition $\theta_{0}$ is in $C^{\infty}_{c}(\mathbb{R}^{d})$, then, for every $t>0$, the solution $\theta_{t}$ is also in $C^{\infty}_{c}(\mathbb{R}^{d})$.
@@
The first assertion is deduced from [10, Proposition 2.1], in the case where $Q$ satisfies Assumption 1.2, or from [16, Theorem 1.3] if $Q$ satisfies Assumption 1.1. In the latter case, the estimate (1.8) follows again from [16, Theorem 1.3], while (1.9) follows similarly by taking into account Remark 3.2 therein. The third assertion follows from the representation formula (2.5) in [10] and the smoothness of the associated stochastic flow (see also the comments before [10, Proposition 2.1]). ∎
If $\theta$ is a solution to (1.1), $\mathbb{P}$-a.s., $\theta$ is in $L^{\infty}_{t}(H^{-d/2-\varepsilon})$ by Sobolev embedding, hence $\Delta\theta$ is in $L^{\infty}_{t}(H^{-d/2-2-\varepsilon})$. By Proposition 1.4 (see also the proof of Proposition 2.6), we have ($\langle\xi\rangle:=(1+|\xi|^{2})^{1/2}$)
@@
Hence $\sum_{k}\|\nabla\cdot(\theta_{s}\sigma_{k})\|_{H^{-d/2-2}}^{2}$ is in $L^{\infty}$ and so
@@
Let $\epsilon>0$, $\theta_{0}\in L^{1}\cap L^{2}$ and let $\theta$ be a weak solution of (1.1), as Definition 1.5, with $\theta_{0}$ as the initial data. Then $\theta$ is $H^{-d/2-\epsilon}$-mild solution to (1.7).
@@
The proof is similar to the proof from [7, Chapter 6]. Let $(\theta_{s})_{s\in[0,T]}$ be a weak solution, in the sense of Definition 1.5. Using a density argument one can show that for all $f\in C^{1}([0,t];\mathcal{S}(\mathbb{R}^{d}))$ we have
$\langle\theta_{t},f_{t}\rangle-\langle\theta_{0},f_{0}\rangle=\int_{0}^{t}\langle\theta_{s},\nabla f_{s}\cdot V(\,\mathrm{d}s)\rangle+\int_{0}^{t}\big{\langle}\theta_{s},\dot{f}_{s}+(\kappa+\nu)\Delta f_{s}\big{\rangle}\,\mathrm{d}s.$
We choose $f_{s}=P_{t-s}\phi$, where $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$. Since $\dot{f}_{s}+(\kappa+\nu)\Delta f_{s}=0$, we get
@@
is in $H^{-d/2-\epsilon}$, when $\{\theta_{s}\}_{s\geqslant 0}$ is a predictable process such that
@@
The latter is true by (1.8), when $Q$ satisfies Assumption 1.1. If $Q$ satisfies Assumption 1.2 this bound is implied by Lemma 2.4.
@@
Finally, we point out that, in the case where $Q$ satisfies Assumption 1.2, (1.7) has a space-time stationary solution [10, Proposition 3.1], which we denote by $\Psi(t,x)$. Moreover, from [10, Corollary 3.2], $\mathbb{E}[\Psi(t,x)]=1$, and $\mathbb{E}[\Psi(t,x)^{2}]<\infty$. The correlation function of this field will appear in the statement of our main result. As such, define $V_{\rm eff}$ to be the symmetric matrix, such that
(1.11) $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}\mathbb{E}[\Psi(0,0)\Psi(0,x)]Q(z)\,\mathrm{d}z.$
Observe that, since $Q\in L^{1}$ and $\Psi(t,x)$ has a finite second moment, this integral is finite.

### v2

@@ Abstract.
We consider a diffusion in a Gaussian random environment that is white in time and study the large-scale behavior of the quenched density with respect to the Lebesgue measure. We show that under diffusive rescaling, the fluctuations of the density converge to a Gaussian limit, described by an additive stochastic heat equation. In the case where the environment is divergence-free our result can be interpreted as computing the scaling limit of the first-order correction to the quenched Central Limit Theorem.
@@ 1. Introduction
for an appropriate function $Q:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$. Here $\delta$ means the Dirac delta function.
@@
(1.3) $\mathrm{d}X_{t}=V(t,X_{t})\,\mathrm{d}t+\sqrt{2\kappa}\,\mathrm{d}W_{t},$
where $W$ is a Brownian motion independent of $V$. Then, for almost every realization of $V$, $\theta(t,x)$ is the density of $X_{t}$ with respect to the Lebesgue measure.
@@
As such, studying the large-scale behavior of (1.1), yields ‘local’ information for a diffusion in a (white-in-time) random environment. This can be done in different scaling regimes. Specifically, [1] distinguishes three different scaling regimes: the diffusive regime, the moderate deviation regime, and the large deviation regime. These correspond to studying the diffusion (1.3) under a diffusive rescaling, under a moderate tilting of the diffusion, or the large deviation behavior, respectively. Moreover, in [1], the authors point out a very interesting connection to the KPZ equation and the KPZ universality class [8]. These conjectures were recently proved rigorously in [34, 14] in $d=1$, and for the moderate deviation regime [^1] The model studied in [34, 14] is a discrete analog of (1.1). However, their methods also apply to the continuous case, as is pointed out in [14], Section 6.3.. We also refer to [10] for a study of a related, integrable model, under the large deviation regime.
Here, we are interested in the behavior of (1.1) in the diffusive scaling regime. In [1], it is predicted that, in this regime, the fluctuations of $\theta(t,x)$, viewed as a random field, fall into the Edwards-Wilkinson universality class. To this end, we point out again the reference [17]. There, the authors studied the point-wise behavior of $\theta(t,x)$ and proved that
@@
as $n\rightarrow\infty$, where $q_{t}(x)$ is the standard $d$-dimensional heat kernel with a specific diffusivity and $\Psi(t,x)$ is an appropriate space-time stationary random field (see [17, Section 3] for more details). Instead, what we are interested in is the behavior of
(1.4) $\mathcal{X}_{n}(t,x):=n^{d/2}(\theta^{n}(t,x)-\mathbb{E}[\theta^{n}(t,x)]),$
where $\theta^{n}(t,x):=n^{d}\theta(n^{2}t,nx)$, viewed as a random element of a Sobolev space with negative order. Our main result shows that $\mathcal{X}_{n}(t,x)$ converges in distribution to an explicit Gaussian limit, confirming the predictions of [1], see Theorem 1.9 for the precise statement. We also mention [3, 4, 35] for related results on discrete models.
@@
We end this introduction by describing the structure of the rest of the paper. In Section 1.2 we set up the solution theory for (1.1) and state our assumptions. We present our main result and briefly describe our methods in Section 1.3. In Section 2.1 we collect estimates on the correlation functions of the model and prove some a priori estimates for (1.1). In Section 2.2 we prove our main result, while in Section 2.3 we show that our result still holds under weaker assumptions on the correlation function of the noise, if we assume that the latter is divergence-free.
@@ 1.1. Notation
  • For $x\in\mathbb{R}^{d}$, we write $x^{\ast}$ as the transposition of $x$. The notation $|x|$ stands for the usual Euclidean norm while $\langle x\rangle=(1+|x|^{2})^{1/2}$.
  • Let $\mathcal{S}$ be the space of Schwartz test functions and denote the Fourier transform of a function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$ as
@@
  • For a function $f:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$, we will use interchangeably the notation $f_{t}$ and $f(t)$ to denote the map $x\rightarrow f(t,x)$.
  • Given $T>0$ and $p,q\geq 1$, we denote $L^{p}_{t}L^{q}_{x}$ for the time-dependent space $L^{p}([0,T],L^{q}(\mathbb{R}^{d}))$; similarly, $C_{t}L^{q}_{x}$ and $C_{t}H^{\alpha}_{x}$ are abbreviations of $C([0,T],L^{q}(\mathbb{R}^{d}))$ and $C([0,T],H^{\alpha}(\mathbb{R}^{d}))$, respectively. Sometimes, we replace the subscript $t$ by $T$ to stress the length of the time interval $[0,T]$.
  • We write $a\lesssim b$ to mean that there is some unimportant constant $C>0$ such that $a\leqslant Cb$; to emphasize the dependence of $C$ on some parameters $d,\kappa$, we use the notation $a\lesssim_{d,\kappa}b$.
  • We will denote by $q_{t}(x)=\frac{1}{(2\pi t)^{d/2}}e^{-\frac{|x|^{2}}{2t}}$ the standard $d$-dimensional heat kernel; the notation $I_{d}$ stands for the $d\times d$ unit matrix.
  • We will make use of the notation $x_{1:p}=(x_{1},\dots,x_{p})$ and for a function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we define $g^{\otimes p}:\mathbb{R}^{pd}\rightarrow\mathbb{R}$, $g^{\otimes p}(x_{1:p}):=g(x_{1})g(x_{2})\dots g(x_{p})$.
@@ 1.2. The Setup and Assumptions
(1.5) $\widehat{Q}(\xi)=g(\xi)\bigg(I_{d}-\frac{\xi\xi^{\ast}}{|\xi|^{2}}\bigg),$
where $g(\xi)=g(|\xi|)$ is a continuous nonnegative radial function satisfying $g\in(L^{1}\cap L^{\infty})(\mathbb{R}^{d})$. It is easy to show that $Q(0)=2\nu I_{d}$ for some $\nu>0$.
Observe that the matrix appearing in the right-hand side of (1.5) is the projection to the subspace orthogonal to $\xi$, so that $Q$ is indeed divergence-free. From (1.5) one can deduce that $Q(x)$ is a symmetric matrix for any $x\in\mathbb{R}^{d}$. As $Q$ is a covariance function, it holds $Q(-x)=Q(x)^{\ast}=Q(x)$, i.e. $Q$ is even.
@@
The matrix $Q$ is smooth, even and compactly supported, such that $Q(0)=2\nu I_{d}$.
@@
It can be proved that the space $\mathcal{H}$ consists of continuous, bounded vector fields. We refer to [24, Lemma 2.2] and the discussion below for more details. We also record the following lemma from the same paper, see Lemma $2.3$ therein [^3] Strictly speaking, in [24] the authors consider only the case where Assumption 1.1 holds, but it is easy to see that similar arguments can be used to establish the same results under Assumption 1.2.:
@@
$Q(x-y)=\sum_{n\in\mathbb{N}}\sigma_{k}(x)\sigma_{k}(y)^{\ast},$
@@
$\sum_{k\in\mathbb{N}}\widehat{\sigma}_{k}(\xi)\overline{\widehat{\sigma}_{k}(\eta)^{\ast}}=\widehat{Q}(\xi)\delta(\xi-\eta),\quad\xi,\eta\in\mathbb{R}^{d},$
where the overline means complex conjugate, and $\delta$ is the Dirac delta.
@@
$\lim_{N\to\infty}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\phi(\xi)^{\ast}\Bigg(\sum_{k=0}^{N}\widehat{\sigma}_{k}(\xi)\overline{\widehat{\sigma}_{k}(\eta)^{\ast}}\Bigg)\overline{\psi(\eta)}~\,\mathrm{d}\xi\,\mathrm{d}\eta$
$=\lim_{N\to\infty}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\phi}(x)^{\ast}\Bigg(\sum_{k=0}^{N}\sigma_{k}(x)\sigma_{k}(y)^{\ast}\Bigg)\overline{\widehat{\psi}(y)}~\,\mathrm{d}x\,\mathrm{d}y$
$=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\phi}(x)^{\ast}\,Q(x-y)\,\overline{\widehat{\psi}(y)}~\,\mathrm{d}x\,\mathrm{d}y$
$=\int_{\mathbb{R}^{d}}\phi(\xi)^{\ast}\widehat{Q}(\xi)\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\overline{\operatorname{e}^{{\rm i}y\cdot\xi}\widehat{\psi}(y)}~\,\mathrm{d}y\,\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}\phi(\xi)^{\ast}\widehat{Q}(\xi)\overline{\psi(\xi)}~\,\mathrm{d}\xi,$
where in the second identity we used the fact that $\sum_{k=0}^{N}\sigma_{k}(x)\sigma_{k}(y)^{\ast}\to Q(x-y)$ uniformly on any compact sets. ∎
@@
and, using the fact that $Q(x)$ is an even function, can be written (at least formally) as Itô integral plus correction, that is
@@
where $\nu$ is as in Assumptions 1.1, 1.2. This leads us to the following notion of solution to (1.1) (see also [24, Definition 2.16]):
@@
Let $(\Omega,\mathcal{A},(\mathcal{F}_{t})_{t},\mathbb{P})$ be a given filtered probability space satisfying the standard assumptions, let $V$ be as above. Let $p\in(1,\infty)$ and let $\theta_{0}\in L^{1}\cap L^{p}$. A solution to (1.1) is an $(\mathcal{F}_{t})_{t}$-progressively measurable process $\theta:[0,T]\times\Omega\to L^{1}\cap L^{p}$, satisfying
@@
    $\int_{0}^{t}\langle\theta_{s},\nabla\phi\cdot V(\,\mathrm{d}s)\rangle=\sum_{k}\int_{0}^{t}\langle\nabla\phi,\sigma_{k}\theta_{s}\rangle\,\mathrm{d}B_{k}(s).$
@@
It can be shown, see for example [24, Appendix B], that, when the spectral intensity $g$ in (1.5) is decaying rapidly at infinity, that is, when $Q$ is sufficiently smooth, the Stratonovich formulation (1.1) makes sense and the equivalence with the Itô formulation (1.7) holds rigorously. Hence, the point (ii) in Definition 1.5 is equivalent to:
@@
In the case $Q$ satisfies Assumption 1.2, if the initial condition $\theta_{0}$ is in $C^{\infty}_{c}(\mathbb{R}^{d})$, then, for every $t>0$, the solution $\theta_{t}$ is also in $C^{\infty}_{b}(\mathbb{R}^{d})$.
@@
In the case where $Q$ satisfies Assumption 1.2, existence and uniqueness, as well as smoothness for $\theta_{0}\in C^{\infty}_{c}$, follow from [17, Proposition 2.1] and [31, Theorem 3.1], the representation formula (2.5) in [17] and the smoothness of the associated stochastic flow $X$ (precisely, $\mathbb{P}$-a.s., $X$ and its space derivatives are jointly continuous in $(t,x)$ and $DX$ is non-singular for every $(t,x)$), see e.g. [29] (see also the comments before [17, Proposition 2.1]). In the case when $Q$ satisfies Assumption 1.1, existence and uniqueness follows from [24, Theorem 1.3]; the estimate (1.8) follows again from [24, Theorem 1.3], while (1.9) follows similarly by taking into account Remark 3.2 therein. ∎
If $\theta$ is a solution to (1.1), $\mathbb{P}$-a.s., $\theta$ is in $L^{\infty}_{t}(H^{-d/2-\varepsilon})$ by Sobolev embedding, hence $\Delta\theta$ is in $L^{\infty}_{t}(H^{-d/2-2-\varepsilon})$. By Proposition 1.4 (see also the proof of Proposition 2.7 below), we have ($\langle\xi\rangle:=(1+|\xi|^{2})^{1/2}$)
@@
Hence $\sum_{k}\|\nabla\cdot(\theta_{s}\sigma_{k})\|_{H^{-d/2-2}}^{2}$ is in $L^{\infty}_{t}$, $\mathbb{P}$-a.s., and so
@@
Let $\epsilon>0$, $\theta_{0}\in L^{1}\cap L^{2}$ and let $\theta$ be a weak solution of (1.1), as in Definition 1.5, with $\theta_{0}$ as the initial data. Then $\theta$ is $H^{-d/2-\epsilon}$-mild solution to (1.7).
@@
The proof is similar to the proof from [13, Chapter 6]. Let $(\theta_{s})_{s\in[0,T]}$ be a weak solution, in the sense of Definition 1.5. Using a density argument one can show that for all $f\in C^{1}([0,t];\mathcal{S}(\mathbb{R}^{d}))$ we have
$\langle\theta_{t},f_{t}\rangle-\langle\theta_{0},f_{0}\rangle=\int_{0}^{t}\langle\theta_{s},\nabla f_{s}\cdot V(\,\mathrm{d}s)\rangle+\int_{0}^{t}\big\langle\theta_{s},\partial_{s}f_{s}+(\kappa+\nu)\Delta f_{s}\big\rangle\,\mathrm{d}s.$
We choose $f_{s}=P_{t-s}\phi$, where $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$. Since $\partial_{s}f_{s}+(\kappa+\nu)\Delta f_{s}=0$, we get
@@
is in $H^{-d/2-\epsilon}$, when $\{\theta_{s}\}_{s\geqslant 0}$ is a progressively measurable process such that
@@
The latter is true by (1.8), when $Q$ satisfies Assumption 1.1. If $Q$ satisfies Assumption 1.2 this bound can be obtained by adapting the proof of Lemma 2.4 (see also Remark 2.5).
@@
Finally, we point out that, in the case where $Q$ satisfies Assumption 1.2, [17, Proposition 3.1 and Corollary 3.2] construct a space-time stationary solution [^4] By stationary field we mean a field whose law is invariant under space-time translations $f\mapsto f(\cdot+s,\cdot+y)$. Technically, to define $\Psi$ as solution to (1.1) in our setting, we should extend Definition 1.5 to include negative times and data in $L^{1}_{loc}$. While this is possible (as in [17]), in this paper we only need that $\Psi$ satisfies (1.16) and (2.24). to (1.1), which we denote by $\Psi(t,x)$, such that $\mathbb{E}[\Psi(t,x)]=1$, and $\mathbb{E}[\Psi(t,x)^{2}]<\infty$. The correlation function of this field will appear in the statement of our main result. As such, define $V_{\rm eff}$ to be the symmetric matrix, such that
(1.11) $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}\mathbb{E}[\Psi(0,0)\Psi(0,z)]Q(z)\,\mathrm{d}z.$
Observe that, since $Q\in L^{1}$ and $\Psi(t,z)$ has a finite second moment, this integral is finite.